Circular No. 7831

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002aw IN ANONYMOUS GALAXY T. Puckett and M. Peoples, Mountain Town, GA, report the discovery

of an apparent supernova (mag 16.6) on an unfiltered CCD frames (limiting mag 20.0) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Feb. 15.39 and 16.47 UT. SN 2002aw is located at α = 16^{h}37^{m}29^{s}.06, δ = +40^{o}52′50′′.3 (equinox 2000.0), which is 1′′.6 west and 1′′.5 north of the center of its apparent host galaxy. S. Nakano, Sumoto, Japan, reports that a confirming unfiltered CCD image of SN 2002aw (mag ∼ 15.5, includes part of galaxy) was also obtained by R. Kushida, Yatsugatake South Base Observatory, on Feb. 16.75; Y. Kushida provides the following position end figures: 29^{s}.06, 49′′.0. The new object was not present on nine images taken by Puckett from 2000 June 29 to 2001 Sept. 17 (limiting mag ∼ 20.0) or on Palomar Sky Survey images taken on 1993 Apr. 30, 1989 May 31 (limiting mag ∼ 21.0), and 1953 May 19 (limiting mag ∼ 19.7).

AQUILA X-1 J. H. Swank and E. A. Smith, Goddard Space Flight Center, report:

“Rossi X-Ray Timing Explorer observations indicate that Aquila X-1 has begun an x-ray outburst. The ASM reported a level of 33 mCrab during Feb. 12–15. On Feb. 15.6 UT, the flux began a rise at the rate of 8.3 ASM counts/s (122 mCrab) per day. The PCA watched as it rose from 75 to 120 mCrab during Feb. 16.03–16.25. One x-ray burst was observed at the level of 1 Crab. Additional RXTE observations are scheduled.”

V838 MONOCEROTIS H. U. Käufl, European Southern Observatory (ESO), reports that ob-

servations by G. Locurto, F. Kerber, B. Heijligers, and himself with the ESO 3.6-m telescope (+ TIMMI2) yield the following infrared magnitudes: Jan. 27.219 UT, [11.9 μm] = 4.38; 28.183, [11.9 μm] = 4.51; 28.195, [8.9 μm] = 4.40; 28.217, [10.4 μm] = 4.70; Feb. 9.125, [11.9 μm] = 2.65; 11.131, [11.9 μm] = 2.93; 11.140, [10.4 μm] = 3.15. A low-resolution spectrum (range 8.0–13.3 μm; λ/Δλ ∼ 150) taken on Feb. 9.167 shows a blackbody continuum corresponding to T ∼ 4500 K and the following lines: 11.3 μm, PAH; 11.6 μm, –; 12.8 μm, [Ne II]. A similar spectrum obtained on Feb. 11.104 shows a 10–15-percent broad silicate absorption feature not present in the previous spectrum, indicating the onset of dust formation around the object, which may lead rapidly to high extinction at shorter wavelengths (as in V4334 Sgr).

2002 February 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green